How many moles of NH3 are there in 77.5 g of NH3?
Answer #1

mass of NH3   = 77.5g

gram molar mass of NH3 = 14+3 = 17g/mole

no of moles of NH3   = mass of NH3/gram molar mass of NH3

                                 = 77.5/17   = 4.56moles

no of moles of NH3 = 4.56moles >>>>answer